Write an equation that you could solve using these steps:
Add 7 to both sides. Then, divide both sides by 4.
Explain how you know your answer is right.

Answer by checkley79(3341)

4x-7=9 add 7 to both sides.
4x=9+7
4x=16 divide by 4
x=16/4
x=4 ans.